You prepared a phosphate buffer solution by mixing
Answer #1

40 mL of 0.25 M KH2PO4 and 60 mL of 0.45 M K2HPO4

total volume is 100 mL which contains

40 mL x 0.25 moles/L of KH2PO4 = (40 x 0.25)/1000 moles = 0.01 moles in 100 mL = (0.01 x 1000)/100 = 0.1 M KH2PO4 in the final solution

60 mL x 0.45 moles/L of K2HPO4 = (60 x 0.45)/1000 moles = 0.027 moles in 100 mL = (0.027 x 1000)/100 = 0.27 M K2HPO4 in the final solution

Use the Henderson Hasselbalch equation

pH = pKa + log [A- ] / [HA]

pH = 7.20 + log (0.27) / (0.1)

pH = 7.20 + log 2.7

pH = 7.63

HCl will react with HPO42- to produce H2PO4-

net ionic equation H+(aq) + HPO42-(aq) \rightarrow H2PO4-(aq)

The added HCl will react with K2HPO4 as shown above

so the concentration of K2HPO4 will decrease and KH2PO4 will increase

the new concentration of K2HPO4 will be 0.027 moles - 0.0060 mol = 0.021 mole in 100 mL = 0.21 M

The new concentration of KH2PO4 will be 0.01 moles + 0.0060 mol = 0.016 mole in 100 mL = 0.16M

The pH will be

pH = pKa + log [A- ] / [HA]

pH = 7.20 + log (0.21) / (0.16)

pH = 7.20 + log 1.31

pH = 7.32